OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 59 miles at a certain speed. The canoeist travels 17 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 5 hrs. What was the speed on each part of the trip?

OpenStudy (anonymous):

round to the nearest hundredth

OpenStudy (anonymous):

okay, we will do this one together

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we have to use the formula, d=r*t

OpenStudy (anonymous):

for the first part of the trip: a canoeist travels 59 miles at a certain speed, which means that we can now write our equation d=r*t as 59=(r*t) sub 1. The reason i right sub1 is becuase we will have to equation like this one and i want to be able to distinguish between both okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now, The canoeist travels 17 miles on the second part of the trip at a speed 5 mph slower. This means that our equation d=r*t, can be written as 17=((r-5)*t)sub 2, okay.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the reason i wrote r-5 is because the canoest is traveling 5 miles slower

OpenStudy (anonymous):

to recap, we have: 59=(r*t) sub 1 and 17=((r-5)*t)sub 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now, the total time for the trip is 5hrs, which means that our last equation can be written as: t sub1+t sub 2=3 okay

OpenStudy (anonymous):

t represnets the time , so logically, the two t variables in our equations when added together give the total time of 3 hrs

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we have 3 equations: 59=(r*t) sub 1 17=((r-5)*t)sub 2 and t sub1+t sub 2=3 okay

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lets begin with the first equation, we have 59=(r*t) sub 1, now we can rearrgane this okay, that way we can solve for tsub1, now in this case tsub1=59/r 17=((r-5)*t)sub 2, same this with the second equation: tsub2=17/r-5

OpenStudy (anonymous):

now with these to values, we look a the third equations: t sub1+t sub 2=3. We now know what tsub 1 and tsub2 represent. We simply plug in those values. and we get (17/r-5)+ (59/r)=3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Still stuck

OpenStudy (anonymous):

made a mistake on it should be: (17/r-5)+ (59/r)=5

OpenStudy (anonymous):

ohhh okay so now I am way confused

OpenStudy (anonymous):

now we will take that new equation, and multiply it through by the common denominator of r(r-5). It will look like this: (r(r-5))((17/r-5)+ (59/r)=3), which after evluation will be: 59(r-5)+17(r)=5(r(r-5))

OpenStudy (anonymous):

what numbers would I plug in for r?

OpenStudy (anonymous):

now all we do is distribute: 59r-295+17r=5r^2-25r

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we simply collect like terms, and we get: 5r^2-101r+295

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

that is one of the last steps.

OpenStudy (anonymous):

then either use the quadratic formula to solve this, or a calculator. I get x as being: x=3.5417 or x=16.568 since we cannot subtract 5 from 3.5417 (we would get a negative value) the we take 16.568 as being our answer. Which means that the speed on the first part of the trip was about 16.568mph